# Improved Off-policy Reinforcement Learning in Biological Sequence Design

### v1

@@ 6.4 Anti-microbial peptide design
Task setup. The goal is to generate protein sequences with anti-microbial properties (AMP). The vocabulary size $|\mathcal{V}|=20$, and the sequence length ($L$) varies across sequences, and we consider sequences of length 50 or lower. For the AMP task, we consider a much larger query batch size for each active round because they can be efficiently synthesized and evaluated (Jain et al., ). We set $\delta$ as 0.5 with $\lambda=1$.
(a) Proxy vs. true scores at $t=1$
(b) $\rho(f,f_{\phi})$
Figure 4: Proxy failure on Hard TF-Bind-8. (a) shows the proxy values (i.e., reward) and true score on the whole data point at the initial round. In (b), the correlation between $f$ and $f_{\phi}$ is much higher when the data points are close to the initial dataset (‘$\leq 0$’ and ‘$\leq 8$’ correspond to the initial dataset and the whole sequence space, respectively).
@@
Results. The results in Table 2 illustrate that ours consistently gives improved performance over GFN-AL regardless of acquisition function. According to the work from Jain et al. , we also adopt conservative model-based optimization method, (COMs; Trabucco et al., 2021) and on-policy reinforcement learning, DyNA PPO (Angermueller et al., 2020) as baselines. Our method demonstrated significantly higher performance in terms of mean, diversity, and novelty compared to the baselines.
6.5 Achieving Pareto frontier with balancing capability of $\delta$-CS
In this analysis, we demonstrate that $\delta$-CS achieves a balanced search using $\delta$, producing Pareto frontiers or comparable results to the baseline methods: GFN-AL (Jain et al., ) and GFNSeqEditor (Ghari et al., 2023). Notably, GFN-AL can be seen as a variant of our method with $\delta=1$, which fully utilizes the entire trajectory search. This approach is expected to yield high novelty and diversity, but it is also prone to generating high rewards due to the increased risk of out-of-distribution samples affecting the proxy model. GFNSeqEditor, on the other hand, leverages GFlowNets as a prior, editing from a wild-type sequence. It is designed to deliver reliable performance and be more robust to out-of-distribution issues by constraining the search to sequences similar to the wild type. However, unlike $\delta$-CS, GFNSeqEditor does not utilize such obtained samples for training GFlowNets in full trajectory level; GFNSeqEditor is expected to have lower diversity and novelty compared to GFN-AL and $\delta$-CS.
As shown in Fig. 3, GFN-AL generally produces higher diversity and novelty in the RNA and GFP tasks compared to GFNSeqEditor. However, GFNSeqEditor performs better in terms of reward on the large-scale GFP task, whereas GFN-AL struggles due to the lack of a constrained search procedure in such a large combinatorial space. In contrast, $\delta$-CS achieves Pareto-optimal performance compared to both methods, clearly outperforming GFNSeqEditor across six tasks, with higher rewards, diversity, and novelty. For the RNA and GFP tasks, we achieve higher scores than GFN-AL while maintaining similar novelty but slightly lower diversity. In the AAV task, $\delta$-CS shows a distinct Pareto improvement. These results demonstrate that $\delta$-CS provides a beneficial balance by combining conservative search with amortized inference on full trajectories using off-policy GFlowNets training, effectively capturing the strengths of both GFN-AL and GFNSeqEditor.
6.6 Study on proxy failure and conservativeness effect
Task: Hard TF-Bind-8. By modifying the initial dataset distribution and the landscape, we can make a harder version of TF-Bind-8. Specifically, we collect the initial dataset near a certain sequence (considered as a wild-type) while ensuring that the initial sequences have lower scores than the given sequence, which is 0.431. The size of $\mathcal{D}_{0}$ is 1,024. Furthermore, we modify the landscape to give 0 rewards for sequences with scores lower than 0.3. These features are often observed in protein design tasks, where the search space is extremely large, e.g., $20^{238}$ previously–with a limited real-world dataset, and the score often falls to 0.
Proxy failure. As shown in Figure 2, $\delta$-CS gives slightly better performance than GFN-AL, but only marginally. This is because the TF-Bind-8 task is relatively easy to optimize, leading to similar results across methods. To more clearly assess the effectiveness of $\delta$-CS, we conduct several studies on the harder TF-Bind-8 task, which is more difficult to optimize. Figure 4(a) illustrates the proxy values and true scores for all $x\in\mathcal{X}$ in the first round. While the proxy provides accurate predictions for the initial data points $x\in\mathcal{D}_{0}$ (represented by the red dots), it produces unreliable predictions for points outside $\mathcal{D}_{0}$. This supports our hypothesis that the proxy model performs poorly on out-of-distribution data.
Effect of $\delta$ conservativeness. Figure 4(b) illustrates that the correlation between the oracle $f$ and the proxy $f_{\phi}$ significantly decreases as data points move farther from the observed dataset. This strongly motivates the use of $\delta$-CS, which constrains the search bounds using $\delta$. By limiting the search to within these constrained edit distances, $\delta$-CS enhances the correlation with the oracle.
Studies on $\delta$. We study on the choice of $\delta$ and effectiveness of adaptive $\delta(x,\sigma)$, in Section B.1.
7 Conclusions
Appendix A Implementation detail
A.1 Proxy training
For training proxy models, we follow the procedure of (Jain et al., ). We use Adam (Kingma, 2015) optimizer with learning rate $1\times 10^{-5}$ and batch size of 256. The maximum proxy update is set as 3000. To prevent over-fitting, we use early stopping using the $10\%$ of the dataset as a validation set and terminate the training procedure if validation loss does not improve for five consecutive iterations.
A.2 Policy training
As described in Section 6, we employ a two-layer long short-term memory (LSTM; Hochreiter & Schmidhuber, 1997) with 512 hidden dimensions. The policy is trained with a learning rate of $5\times 10^{-4}$ with a batch size of 256. The learning rate of $Z$ is set as $10^{-3}$. The coefficient $\kappa$ in Equation 2 is set as 0.1 for TF-Bind-8 and AMP with MC dropout, according to Jain et al. , and 1.0 for RNA and protein design with Ensemble following Ren et al. (2022).
A.3 Implementation details of baselines
We adopt open-source code from FLEXS benchmark (Sinai et al., 2020).
  • GFN-AL (Jain et al., ): We strictly follow hyperparameters of the original code in they conduct experiments on TF-Bind-8 and AMP. The proxy is parameterized using an MLP with two layers of 2,048 hidden. For the policy, a 2-layer MLP with 2,048 hidden dimensions is used, but we also test it with a 2-layer LSTM policy.
Appendix B Further studies
B.1 Studies on effect of $\delta$ on Hard TF-Bind-8
(a) Median scores with various fixed $\delta$
(b) Effect of adaptive control
@@ B.1 Studies on effect of $\delta$ on Hard TF-Bind-8
To verify its effectiveness and give intuition about how to set $\delta$, we conduct experiments with various $\delta$ in Hard TF-Bind-8. The results show that $\delta$-CS with $\delta<1$ can significantly outperform GFN-AL by searching for data points that correlate better with the oracle. In the Hard TF-Bind-8 task, a more conservative search with $\delta=0.25$ is beneficial since the proxy is unreliable in the early rounds. In particular, ours with $\delta=1$ means the full on-policy search (no conservativenss). The performance differences between $\delta=1$ and GFN-AL came from ${\epsilon}$-noisy behavior policy, which selects random actions with a probability of 0.001, in GFN-AL. Furthermore, using adaptive $\delta(x,\sigma)$ mostly gives the improved scores as depicted in Figure 5(b).
We further examine the effect of adaptive $\delta$ adjustment on RNA in the following section.
B.2 Studies on the effect of adaptive $\delta$ on RNA-A
@@
We examine the effects of proxy uncertainty-based $\delta$. In RNA, the average proxy standard deviation $\bar{\sigma}$ at the initial round is observed as 0.005 to 0.012. Therefore, we set $\lambda=5$ to roughly make $\lambda\bar{\sigma}\approx 1/L$, where $L=14$. As illustrated in Figure 6, $\delta(x;\sigma)$ consistently gives the higher score. However, the constant $\delta=5$ still outperforms all baselines, exhibiting the robustness of $\delta$-CS.
B.3 Balancing capability of $\delta$ on RNA
Similar to Section 6.5, we also verify the balancing capability of $\delta$-CS on RNA-B and RNA-C. The $\delta$ is set as 0.5.
Figure 7: Average score and diversity/novelty with five independent runs on RNA design tasks.
B.4 Studies on rank-based reweighted sampling in proxy training
The rank-based reweighing also can be used in proxy training, i.e., $x\sim P_{\mathcal{D}_{t-1}}(x;k)$, where $k$ is a reweighting factor and fixed as 0.01 in this work. The results show that rank-based reweighted proxy training improves performance mostly. However, the gap is small, and $\delta$-CS still works well even without reweighting.
Table 3: Ablation studies of rank-based reweighted proxy training
Max Median Mean Diversity Novelty
RNA-A with rank-based 1.055 $\pm$ 0.000 0.939 $\pm$ 0.008 0.947 $\pm$ 0.009 6.442 $\pm$ 0.525 7.406 $\pm$ 0.066
without rank-based 1.049 $\pm$ 0.010 0.936 $\pm$ 0.016 0.944 $\pm$ 0.015 5.782 $\pm$ 0.697 7.397 $\pm$ 0.098
RNA-B with rank-based 1.014 $\pm$ 0.001 0.929 $\pm$ 0.004 0.934 $\pm$ 0.003 5.644 $\pm$ 0.307 7.661 $\pm$ 0.064
without rank-based 1.009 $\pm$ 0.008 0.932 $\pm$ 0.012 0.938 $\pm$ 0.012 6.252 $\pm$ 0.291 7.673 $\pm$ 0.033
RNA-C with rank-based 1.094 $\pm$ 0.045 0.972 $\pm$ 0.043 0.983 $\pm$ 0.043 6.493 $\pm$ 1.751 6.494 $\pm$ 0.084
without rank-based 1.097 $\pm$ 0.022 0.958 $\pm$ 0.029 0.965 $\pm$ 0.031 5.472 $\pm$ 1.921 6.464 $\pm$ 0.192

### v2

@@ B.1 Anti-microbial peptide design
Results. The results in Table 1 illustrate that ours consistently gives improved performance over GFN-AL regardless of acquisition function. According to the work from Jain et al. (2022), we also adopt conservative model-based optimization method, (COMs; Trabucco et al., 2021) and on-policy reinforcement learning, DyNA PPO (Angermueller et al., 2020) as baselines. Our method demonstrated significantly higher performance in terms of mean, diversity, and novelty compared to the baselines.
B.2 $\delta$-CS with various off-policy RL
We extend our experiments on protein design tasks to include three soft off-policy RL algorithms: Soft Q-Learning (SQL) (Haarnoja et al., 2017), Log-Partition Variance Gradient (VarGrad; Richter et al., 2020), and Soft Path Consistency Learning (Soft PCL) (Nachum et al., 2017; Chow et al., 2018; Madan et al., 2023), yielding four baselines in total (TB, SQL, VarGrad, and Soft PCL). Trajectory Balance (TB) and VarGrad apply constraints over full trajectories, resulting in higher variance but lower bias. TB explicitly estimates state flows or terminal partition functions, whereas VarGrad does so implicitly via batch averaging. Soft PCL operates over sub-trajectories (akin to TD-$\lambda$), balancing bias and variance through trajectory length. SQL relies on one-step transitions, reducing variance but increasing bias. For a broader discussion on connections between off-policy RL in discrete sampling, see the works by Tiapkin et al. (2024) and Deleu et al. (2024).
Table 2: Mean score of Top-128 after ten active rounds
GFP AAV
SQL 3.331 $\pm$ 0.000 0.480 $\pm$ 0.000
SQL + $\delta$-CS 3.573 $\pm$ 0.003 0.495 $\pm$ 0.001
VarGrad 3.331 $\pm$ 0.000 0.480 $\pm$ 0.000
VarGrad + $\delta$-CS 3.567 $\pm$ 0.003 0.668 $\pm$ 0.011
Soft PCL ($\lambda$=0.9) 3.331 $\pm$ 0.000 0.480 $\pm$ 0.000
Soft PCL ($\lambda$=0.9) + $\delta$-CS 3.578 $\pm$ 0.004 0.622 $\pm$ 0.011
As shown in Table 2, $\delta$-CS consistently improves performance across all methods, demonstrating its plug-and-play versatility. Since the global sequence structure is crucial in protein design, full-trajectory methods (TB and VarGrad with $\delta$-CS) tend to outperform shorter-horizon approaches.
B.3 Comparison with back-and-forth search in LS-GFN
Our search with noise injection and denoising policies shares similar insight with a back-and-forth search in local search GFlowNets (LS-GFN; Kim et al., 2024d). They both induce new sequences by partially revising given sequences. In particular, LS-GFN incorporates a back-and-forth search strategy that partially backtracks trajectories using a backward policy and reconstructs them using a forward policy of the GFN. This section compares the search strategies under the conservative search; we control the conservativeness of LS-GFN by setting the backtracking step $K$ as $\delta L$.
(a) RNA-B
(b) RNA-C
(c) AAV
Figure 7: Comparison $\delta$-CS with back-and-forth search in LS-GFN on RNA-B,C and AAV
B.4 Studies with smaller initial dataset on protein designs
In this section, we extend our ablation study by comparing our method, $\delta$-CS, against two baselines: GFN-AL and an additional off-policy search method, LS-GFN (Kim et al., 2024d). This study evaluates the performance of $\delta$-CS under varying initial dataset sizes, $|\mathcal{D}_{0}|=1,000$, and compares the results to the original ablation study setup.
As shown in Table 3, our $\delta$-CS demonstrates a substantial advantage over both GFN-AL and its improved variant, LS-GFN, which leverages back-and-forth search. The results highlight the effectiveness of $\delta$-CS in enhancing GFN training by enabling a more robust and conservative off-policy search, which is critical for improving proxy-based active learning.
Table 3: Ablation study results with 1,000 initial datapoints for GFP and AAV tasks, showing maximum values achieved after active learning.
Method GFP AAV
Adalead 3.568 $\pm$ 0.005 0.557 $\pm$ 0.023
GFN-AL 3.586 $\pm$ 0.006 0.560 $\pm$ 0.008
GFN-AL + LS (Kim et al., 2024d) 3.580 $\pm$ 0.003 0.493 $\pm$ 0.006
GFN-AL + $\delta$-CS 3.591 $\pm$ 0.007 0.704 $\pm$ 0.024
To further verify that $\delta$-CS is robust to proxy misspecification compared to other GFN methods, we conducted additional experiments to test whether this hypothesis holds at different levels of proxy model quality.
To degrade the proxy model quality, we truncated the initial dataset at different levels—$50\%$, $25\%$, and $10\%$ percentiles based on reward values. Proxy models trained on datasets with lower percentile cutoffs are more misspecified for higher-reward data points, making GFN training and search more challenging. Under these circumstances, we compared our method with GFN-AL and LS-GFN (Kim et al., 2024d) as GFN baselines.
Figure 8: Maximum values achieved after active learning with varying initial dataset quality (AAV task).
As shown in Figure 8, the performance decreases as the percentile decreases, which is expected because the proxy quality deteriorates significantly. Among the baselines, our method consistently provides substantially better performance than the others. This demonstrates that our hypothesis—that a conservative search with $\delta$-CS is necessary—holds across different levels of proxy model quality.
B.5 Experiments on different GFP benchmark
The GFP task has multiple benchmark variants, primarily due to differences in oracle models (e.g., TAPE (Rao et al., 2019), Design-Bench Transformer (Trabucco et al., 2022)) and scoring normalization strategies. Some studies use normalized oracle scores (Kirjner et al., 2024; Lee et al., 2024), while others adopt raw scores (Sinai et al., 2020; Jain et al., 2022; Ren et al., 2022); see the work by  Surana et al. (2024) for a comprehensive comparison.
To enable consistent comparison across benchmarks, we follow the experimental setup of LatProtRL (Lee et al., 2024). Specifically, we run 15 active learning rounds, evaluating 256 sequences per round, using the oracle model and initial dataset split as in the work by Kirjner et al. (2024). Following LatProtRL, we report the median top-128 score across 5 independent runs. Results for AdaLead and LatProtRL are directly obtained from the work by Lee et al. (2024).
Table 4: GFP optimization results on the benchmark from Kirjner et al. (2024) and Lee et al. (2024)
GFP Medium GFP Hard
AdaLead 0.93 $\pm$ 0.0 0.75 $\pm$ 0.1
LatProtRL 0.93 $\pm$ 0.0 0.85 $\pm$ 0.0
GFN-AL ($\delta=1$) 0.21 $\pm$ 0.0 0.09 $\pm$ 0.0
GFN-AL + $\delta$-CS ($\delta=0.05$) 0.57 $\pm$ 0.0 0.60 $\pm$ 0.1
GFN-AL + $\delta$-CS ($\delta=0.01$) 1.06 $\pm$ 0.1 0.86 $\pm$ 0.0
As shown in Table 9, $\delta$-CS consistently outperforms baseline methods across all benchmark variants. In particular, it significantly improves upon GFN-AL (equivalent to $\delta=1$, i.e., without conservative search), achieving the best results with $\delta=0.01$. This benchmark is intentionally designed to be more challenging compared to Design-Bench, as the initial datasets in GFP-medium and GFP-hard consist of low-scoring and distant sequences to the 99th percentile sequences; see Appendix A in the work by Kirjner et al. (2024). This results in greater proxy uncertainty, which in turn necessitates a lower $\delta$ to achieve effective exploration.
B.6 Studies on effect of adaptive $\delta$
This section provides the experimental results to verify the effectiveness of considering uncertainty on each data point for adjusting the conservativeness.
B.6.1 Hard TF-Bind-8
@@
To verify its effectiveness and give intuition about how to set $\delta$, we conduct experiments with various $\delta$ in Hard TF-Bind-8. The results show that $\delta$-CS with $\delta<1$ can significantly outperform GFN-AL by searching for data points that correlate better with the oracle. This holds even when we use constant values for $\delta$. However, as depicted in Figure 9, using adaptive $\delta(x,\sigma)$ mostly gives the improved scores.
B.6.2 RNA sequence design
@@
We examine the effects of proxy uncertainty-based $\delta$. In RNA, the average proxy standard deviation at the initial round is observed as 0.005 to 0.012. Therefore, we set $\lambda=5$ to roughly make $\lambda\bar{\sigma}\approx 1/L$, where $L=14$. As illustrated in Figure 10, $\delta(x;\sigma)$ consistently gives the higher score. However, the constant $\delta=5$ still outperforms all baselines, exhibiting the robustness of $\delta$-CS.
B.7 Sensitivity analysis over various $\delta_{\text{const.}}$
This section provides the experimental results with varying $\delta_{\text{const}}$ on RNA and protein designs. Note that we use $\delta_{\text{const}}=0.5$ in the main experiments, meaning that approximately 7 tokens are masked in RNA designs, while we use $\delta_{\text{const}}=0.05$ for protein designs–approximately masking 4-12 tokens for GFP ($L=238$) and AAV ($L-90$).
(a) RNA-A
(b) RNA-B
(c) RNA-C
Figure 11: Adaptive delta with various $\delta_{\text{const}}$ on RNA
(a) GFP
(b) AAV
Figure 12: Adaptive delta with various $\delta_{\text{const}}$ on protein designs
The results in Figure 11 and Figure 12 show that GFN-AL with $\delta$-CS consistently gives improved results compared to GFN-AL. Even in AAV, $\delta_{\text{const}}=0.075$ outperforms other baseline, including GFN-AL whose score is 0.509. Moreover, we obtain even better results by setting $\delta_{\text{const}}=0.01$ in GFP and $\delta_{\text{const}}=0.025$ in AAV, indicating that domain-specific tuning of $\delta_{\text{const}}$ can yield further gains.
B.8 Balancing capability with various $\delta$
Similar to Section 6.3, we evaluate $\delta$-CS on RNA-B and RNA-C by varying $\delta_{\text{const}}$ from 0.1 to 0.5. As illustrated in Figure 13, $\delta$-CS consistently reaches the Pareto frontier. In the protein designs, we vary $\delta_{\text{const}}$ between 0.01 and 0.05. Figure 14 demonstrates that GFN-AL with $\delta$-CS forms the Pareto curve over both GFN-AL and GFNSeqEditor, further confirming the balancing capability of $\delta$-CS.
Figure 13: Average score and diversity/novelty with on RNA designs with various $\delta$.
Figure 14: Average score and diversity/novelty on protein designs with various $\delta$.